# MigGPT: Harnessing Large Language Models for Automated Migration of Out-of-Tree Linux Kernel Patches Across Versions

### v1

@@ Abstract
Out-of-tree kernel patches are essential for adapting the Linux kernel to new hardware or enabling specific functionalities. Maintaining and updating these patches across different kernel versions demands significant effort from experienced engineers. Large language models (LLMs) have shown remarkable progress across various domains, suggesting their potential for automating out-of-tree kernel patch migration. However, our findings reveal that LLMs, while promising, struggle with incomplete code context understanding and inaccurate migration point identification. In this work, we propose MigGPT, a framework that employs a novel code fingerprint structure to retain code snippet information and incorporates three meticulously designed modules to improve the migration accuracy and efficiency of out-of-tree kernel patches. Furthermore, we establish a robust benchmark using real-world out-of-tree kernel patch projects to evaluate LLM capabilities. Evaluations show that MigGPT significantly outperforms the direct application of vanilla LLMs, achieving an average completion rate of 72.59% ($\uparrow 50.74\%$) for migration tasks.
Machine Learning, ICML
@@ 1 Introduction
The Linux kernel, a widely-used open-source operating system, is extensively applied across various domains (Tan et al., 2020; Lin et al., 2022; de Oliveira et al., 2023). Its adaptability and extensibility enable developers to create out-of-tree kernel patches that enhance performance (Kim et al., 2020; Adam, 2021) or security (Xu et al., 2023; Zhou et al., 2020), contributing to its widespread adoption. Out-of-tree kernel patches, such as RT-PREEMPT, AUFS, HAOC, Raspberry Pi kernel, and Open vSwitch, are modifications to the Linux kernel that are developed and maintained independently of the mainline source tree. Unlike in-tree patches, which are included in official kernel releases, out-of-tree patches address specific use cases or features not yet supported by the mainline kernel. As the Linux kernel evolves, these out-of-tree patches require ongoing maintenance to ensure compatibility with newer Linux kernel versions. As shown in Figure 1, the maintenance process involves utilizing the old out-of-tree kernel patch and analyzing the differences between the old and new Linux kernel versions to upgrade the patched kernel repository to the new version. This maintenance process demands specialized engineers and often takes weeks of intensive effort (Zhang et al., 2021).
@@
Existing code migration technologies (Xing & Stroulia, 2007; Lamothe et al., 2022; Fazzini et al., 2019; Haryono et al., 2020; Thung et al., 2019; Ketkar et al., 2019; Rolim et al., 2017; Dilhara et al., 2023; Shi et al., 2022; Shariffdeen et al., 2021a, b; Yang et al., 2023) utilize static program analysis (Landi, 1992) to facilitate API cross-version maintenance or the backporting of CVE security patches. However, these methods only address a subset of scenarios in out-of-tree kernel patch migration. They rely on predefined migration rules, which are insufficient for handling comprehensive scenarios involving complex changes such as namespace modifications, invocation conflict resolution, and the integration of control and data flow dependencies.
@@
With the substantial progress made by Large Language Models (LLMs) in understanding (Patel et al., 2023; Izadi et al., 2022; Kim et al., 2021) and generating code (Feng et al., 2020; Ahmed & Devanbu, 2022; Ouyang et al., 2023; Yan et al., 2023; Yang et al., 2024), there is a promising opportunity to leverage LLMs for the automated migration and maintenance of out-of-tree kernel patches. However, due to the inherent lack of determinism in LLMs when generating content, several challenges arise when directly employing these models to handle the migration and maintenance of out-of-tree kernel patches. These challenges include 1) interference from similar namespaces, 2) misalignment of code lines, 3) absence of associated code information, and 4) inaccuracies in locating migration points.
@@
To address the challenges , we propose MigGPT, the first framework designed to assist humans in automating the migration and maintenance of out-of-tree kernel patches. MigGPT utilizes Code Fingerprint (CFP), a meticulously designed novel data structure to encapsulate the structural and critical information of code snippets throughout the migration process of out-of-tree kernel patches. With the assistance of CFP, MigGPT incorporates three core modules: the Retrieval Augmentation Module (addressing challenges 1 and challenges 3), the Retrieval Alignment Module (addressing challenge 2), and the Migration Enhancement Module (addressing challenge 4). Specifically, the Retrieval Augmentation Module supplies code snippet information via CFPs, mitigates interference from similar namespaces, and appends additional code snippets pertinent to migration. The Retrieval Alignment Module achieves alignment of the target code snippets through the first anchor line and the last anchor line within CFPs. The Migration Enhancement Module facilitates accurate and efficient migration by comparing CFPs to ascertain the number of migration points and their respective locations.
@@
To evaluate the efficiency of MigGPT, we construct a robust benchmark that includes three real-world projects from the out-of-tree patch community of the Linux kernel. These projects comprise two different levels of migration examples, encompassing a variety of common migration types. With this benchmark, we evaluated MigGPT across diverse LLMs (GPT-3.5, GPT-4-turbo , DeepSeek-V2.5, DeepSeek-V3, and Llama-3.1) (OpenAI, 2023a, b; DeepSeek-AI et al., 2024a, b; Dubey et al., 2024) to validate its effectiveness and broad applicability. MigGPT significantly outperforms the direct application of vanilla LLMs, achieving an average completion rate of 72.59% ($\uparrow 50.74\%$) for migration tasks. Additionally, the average number of queries to LLMs is only 2.22($\uparrow 0.22$), indicating no substantial increase in computational overhead.
@@
  • We have developed a robust and migration benchmark, encompassing three real-world projects. To the best of our knowledge, this is the first benchmark for out-of-tree kernel patch migration that can assess performance across diverse migration tools, providing a valuable foundation for future research.
@@
  • We conduct comprehensive experiments on both closed-source models (i.e. GPT-3.5 and GPT-4) and open-source models (i.e. DeepSeek-V2.5, DeepSeek-V3, and Llama-3.1). The results demonstrate that MigGPT achieved an average migration accuracy of 72.59%($\uparrow 52.74\%$), representing a significant improvement over vanilla LLMs.
@@ 2.1 Code Migration Kernel Patch
Existing code migration (Dilhara et al., 2024; Pan et al., 2024) technologies primarily focus on API cross-version maintenance (Xing & Stroulia, 2007; Lamothe et al., 2022; Fazzini et al., 2019; Haryono et al., 2020; Thung et al., 2019; Ketkar et al., 2019; Rolim et al., 2017; Dilhara et al., 2023) and the backporting of CVE security patches (Shi et al., 2022; Shariffdeen et al., 2021a, b; Yang et al., 2023). Research on API cross-version maintenance employs static analysis (Landi, 1992) to detect deprecated API patterns and migrate code using transformation rules. However, these methods only partially address out-of-tree kernel patch migration due to the tight coupling between kernel and patch code. Similarly, CVE security patch backporting techniques identify vulnerability patterns and apply predefined rules but fail to manage complex changes such as namespace modifications, invocation conflicts, and control/data flow dependencies, limiting their effectiveness in comprehensive migration scenarios.
@@ 2.2 LLMs for Coding
In recent years, LLMs (Chen et al., 2021; Fried et al., 2023; Rozière et al., 2023; Le et al., 2022; Nijkamp et al., 2023; Li et al., 2023; OpenAI, 2023b) have achieved remarkable progress in various natural language processing tasks. Initially focused on natural language understanding and generation, the adaptability of LLMs has expanded to the field of software engineering, where they can be fine-tuned to perform programming tasks such as code completion (Patel et al., 2023; Izadi et al., 2022; Kim et al., 2021), code search (Feng et al., 2020), code summarization (Ahmed & Devanbu, 2022), code generation (Ouyang et al., 2023), and even complex code repair (Fu et al., 2022; Jesse et al., 2023). This inspires us to apply LLMs to the migration and maintenance of out-of-tree kernel patches. To the best of our knowledge, MigGPT is the first work to apply LLMs to this task, paving the way for subsequent research.
3 Preliminaries
Out-of-Tree Kernel Patches
Out-of-tree kernel patches are modifications outside the mainline Linux kernel, often created by third-party developers or organizations to add features, fix issues, or optimize performance for specific hardware. They provide quick access to new functionalities or hardware support, catering to specialized needs. Examples include RT-PREEMPT and AUFS for feature enhancements, HAOC for security improvements, and Raspberry Pi kernel and Open vSwitch for platform-specific optimizations. An example of out-of-tree kernel patch migration is provided in App. B.
Motivation.
Such out-of-tree kernel patches lack official support and require manual maintenance to ensure compatibility with future kernel versions, often demanding weeks of effort from specialized engineers. Automating this process could significantly reduce reliance on manual labor, thereby conserving substantial human resources and accelerating development cycles.
Problem Formulation.
Out-of-tree kernel patches lack official support and require manual maintenance to ensure compatibility with future Linux kernel versions. Let $R$ denote a Linux kernel repository, where $s\in R$ represents a code snippet within the repository. The older version of Linux kernel is $R_{\text{old}}$, and after applying an out-of-tree patch, it becomes $R^{\prime}_{\text{old}}$. When the kernel advances to a new version $R_{\text{new}}$, the migration problem is to construct a function $M:R^{\prime}_{\text{old}}\to R^{\prime}_{\text{new}}$ where $\forall s\in R^{\prime}_{\text{old}},\exists M(s)\in R^{\prime}_{\text{new}}$ s.t.$\forall x\in\text{Inputs},\text{Execute}(R^{\prime}_{\text{old}},x)=\text{Execute}(R^{\prime}_{\text{new}},x)$.
@@ 4.1 Migration Types
Type 1: This type of migration example satisfies $\Delta\neq\varnothing,\Sigma\neq\varnothing$, $\forall\delta\in\Delta,\forall\sigma\in\Sigma,\langle\delta,\sigma\rangle=0$. This indicates that both the out-of-tree kernel patch and the new version of the Linux kernel have modified the code snippet, and their changes do not affect the same lines of code, meaning the modifications do not overlap or conflict with each other.
@@
Table 1: Formalization, Counts of the Two Types of Migration Example. Other examples are too simple to necessitate resolution. More detailed examples are in the App. A.2.
Class Formalization Number Type 1 $\Delta\neq\varnothing,\Sigma\neq\varnothing,$ 80 $\forall\delta\in\Delta,\forall\sigma\in\Sigma,\langle\delta,\sigma\rangle=0$ (59.3%) Type 2 $\Delta\neq\varnothing,\Sigma\neq\varnothing,$ 55 $\forall\delta\in\Delta,\forall\sigma\in\Sigma,\langle\delta,\sigma\rangle\neq 0$ (40.7%) Others $\Delta=\varnothing$ or Too simple $\Delta\neq\varnothing,\Sigma=\varnothing$ to resolve
@@ 4.2 Benchmark Design
We have built a robust migration testing benchmark using out-of-tree kernel patches from real-world projects, specifically focusing on three open-source initiatives: RT-PREEMPT (Kernel, 2013), Raspberry Pi Linux (Pi, 2018), and HAOC (HAOC, 2024)  [^1] Even with knowledge of the code in these out-of-tree kernel patches, LLMs still struggle to accomplish migration and maintenance tasks.. RT-PREEMPT enhances the Linux kernel’s real-time performance for timing-sensitive applications like industrial control and robotics, while Raspberry Pi Linux offers a lightweight kernel optimized for embedded systems. HAOC improves kernel security through a ”dual-kernel” design, enhancing code behavior, data access, and permission management. We collect code from these projects across Linux kernel versions 4.19, 5.4, 5.10, and 6.6 for our benchmark.
@@ 5.1 Challenges
Challenge 1 (Namespace Interference): When identifying the code snippet $s_{\text{new}}$ in $\text{file}_{\text{new}}$ for migration, retrieval errors can occur. LLMs often struggle to locate function definitions within $s_{\text{new}}$ due to interference from similarly named functions, leading to inaccuracies that affect subsequent migration stages. An example is provided in App. C.1.
Challenge 2 (Misalignment): This challenge occurs when retrieving $s_{\text{new}}$ from $\text{file}_{\text{new}}$. Due to the inherent randomness in LLM-generated responses, discrepancies often arise between the start and end lines of $s_{\text{new}}$ identified by the LLM and those retrieved by human developers ($\hat{s}_{\text{new}}$). This misalignment can result in missing or extraneous lines, significantly compromising migration outcomes where precise code segment boundaries are critical. An example is provided in App. C.2.
@@ 5.3 Code Fingerprint
To address the challenges LLMs face in migrating out-of-tree kernel patches across Linux kernel versions, a detailed analysis of code snippet structure is essential to identify migration-related code structure information and code scope constraints. While tools like Abstract Syntax Tree (AST) are useful for structural analysis, they have limitations: 1) Inability to process incomplete or non-compilable code snippets due to tight integration with the compilation process. 2) The mismatch between excessive structural details (AST tools provide a plethora of information irrelevant to patch migration) and the absence of critical information (such as comments and inline assembly), which is essential for maintaining and updating kernel patches  [^2] Inline assembly is widely used in the Linux kernel, and comments are crucial for future module development, as their omission would hinder subsequent modifications.. Focusing on key statements, such as migration points and alignment positions, while preserving essential elements like comments and inline assembly, can enhance efficiency and reduce overhead in the migration process.
To address the limitations of traditional code structure analysis, we propose Code Fingerprint (CFP), a lightweight sequential data structure for analyzing code snippets. CFP records both the content and positional information for each line of statements, encompassing all C language statements, including comments and inline assembly (a detailed example is provided in App. D.1). As shown in Figure 4, CFP focuses on recording function definitions and function calls, which are crucial for addressing challenges 1 and 3, as detailed in Section 5.4. Additionally, its linear structure facilitates accurate positioning for insertion, deletion, and other update operations, tackling challenges 2 and 4, further explained in Sections 5.5 and  5.6. Overall, CFP offers three key advantages: 1) effective processing of incomplete code snippets, 2) preservation of critical information such as comments and inline assembly, which are vital for out-of-tree kernel patch migration, and 3) a streamlined design that focuses on essential statements, improving migration efficiency and reducing overhead. By minimizing unnecessary processing while ensuring relevance, CFP provides a targeted solution for migrating out-of-tree kernel patches across Linux kernel versions.
Figure 3: A C code snippet containing inline assembly statements and comment annotations.

### v2

@@ Abstract
Out-of-tree kernel patches are essential for adapting the Linux kernel to new hardware or enabling specific functionalities. Maintaining and updating these patches across different kernel versions demands significant effort from experienced engineers. Large language models (LLMs) have shown remarkable progress across various domains, suggesting their potential for automating out-of-tree kernel patch migration. However, our findings reveal that LLMs, while promising, struggle with incomplete code context understanding and inaccurate migration point identification. In this work, we propose MigGPT, a framework that employs a novel code fingerprint structure to retain code snippet information and incorporates three meticulously designed modules to improve the migration accuracy and efficiency of out-of-tree kernel patches. Furthermore, we establish a robust benchmark using real-world out-of-tree kernel patch projects to evaluate LLM capabilities. Evaluations show that MigGPT significantly outperforms the direct application of vanilla LLMs, achieving an average completion rate of 74.07% ($\uparrow 45.92\%$) for migration tasks.
@@ 1 Introduction
The Linux kernel, a widely-used open-source operating system, is extensively applied across various domains intro-1 ; intro-2 ; intro-3 ; add1 . Its adaptability and extensibility enable developers to create out-of-tree kernel patches that enhance performance intro-4 ; intro-5 or security intro-6 ; intro-7 ; add2 , holding irreplaceable significance in modern engineering practices. Out-of-tree kernel patches, such as RT-PREEMPT, AUFS, HAOC, Raspberry Pi kernel, and Open vSwitch, are modifications to the Linux kernel that are developed and maintained independently of the mainline source tree. Unlike in-tree patches, which are included in official kernel releases, out-of-tree patches address specific use cases or features not yet supported by the mainline kernel. As the Linux kernel evolves, these out-of-tree patches require ongoing maintenance to ensure compatibility with newer Linux kernel versions. As shown in Figure 1, the maintenance process involves utilizing the old out-of-tree kernel patch and analyzing the differences between the old and new Linux kernel versions to upgrade the patched kernel repository to the new version. This maintenance process is crucial and labor-intensive in engineering applications, which demands specialized experts and takes weeks of intensive effort work .
Existing code migration technologies diff_catchup ; a3 ; app_evovle ; cocci_evolve ; neat ; T2R ; REFAZER ; PYEVOLVE ; SKYPORT ; FixMorph ; PatchWeave ; TSBPORT utilize static program analysis static_program_analysis to facilitate API cross-version maintenance or the backporting of CVE security patches. However, these methods only address a subset of scenarios in out-of-tree kernel patch migration. 1) They rely on predefined migration rules, which are insufficient for handling comprehensive scenarios involving complex changes such as namespace modifications, invocation conflict resolution, and the integration of control and data flow dependencies.
@@
2) These single-step methods assume known target code locations in updated repositories, limiting their applicability to out-of-tree kernel patch scenarios.
@@
With the substantial progress made by Large Language Models (LLMs) in understanding rf-1 ; rf-2 ; rf-3 and generating code codebert ; rf-4 ; rf-5 ; intro-d-1 ; intro-d-2 , there is a promising opportunity to leverage LLMs for the automated migration and maintenance of out-of-tree kernel patches. However, due to the inherent lack of determinism in LLMs when generating content, several challenges arise when directly employing these models to handle the migration and maintenance of out-of-tree kernel patches. These challenges include 1) structural similarity-induced failure, 2) non-deterministic code snippet boundaries, 3) absence of associated code snippet information, and 4) inaccuracies in locating migration points, which reveals that LLMs struggle with incomplete code context understanding and inaccurate migration point identification.
@@
To address the challenges, we propose MigGPT, the first framework designed to assist humans in automating the migration and maintenance of out-of-tree kernel patches. We reframe this migration process as a two-step task: retrieving target code from updated kernels and performing patch migration, which is both more challenging and practical. MigGPT utilizes Code Fingerprint (CFP), a novel data structure to encapsulate the structural and critical information of code snippets throughout the migration process of out-of-tree kernel patches. With the assistance of CFP, MigGPT incorporates three core modules: the Retrieval Augmentation Module (addressing challenges 1 and challenges 3), the Retrieval Alignment Module (addressing challenge 2), and the Migration Enhancement Module (addressing challenge 4). Specifically, the Retrieval Augmentation Module supplies code snippet information via CFPs, mitigates interference from similar structures, and appends additional code snippet information pertinent to migration. The Retrieval Alignment Module achieves alignment of the target code snippet boundaries through the first anchor line and the last anchor line within CFPs. The Migration Enhancement Module facilitates accurate and efficient migration by comparing CFPs to ascertain the number of migration points and their respective locations.
@@
To evaluate the efficiency of MigGPT, we construct a robust benchmark that includes three real-world projects from the out-of-tree patch community of the Linux kernel. These projects comprise two different levels of migration examples, encompassing a variety of common migration types. With this benchmark, we evaluated MigGPT across diverse LLMs (GPT-3.5, GPT-4-turbo, OpenAI-o1, DeepSeek-V3, DeepSeek-R1, and Llama-3.1) gpt-3.5 ; gpt4 ; openai-o1 ; deepseek-3 ; deepseek-r1 ; llama-3.1 to validate its effectiveness and broad applicability. MigGPT significantly outperforms the direct application of vanilla LLMs, achieving an average completion rate of 74.07% ($\uparrow 45.92\%$) for migration tasks. Additionally, the average number of queries to LLMs is only 2.26($\uparrow 0.26$), indicating no substantial increase in computational overhead. Meanwhile, MigGPT required only 2.08% of the average time taken by human experts, demonstrating its superior time efficiency.
@@
  • We have developed a robust migration benchmark, encompassing three real-world projects. To the best of our knowledge, this is the first benchmark for out-of-tree kernel patch migration that can assess performance across diverse migration tools, providing a valuable foundation for future research.
@@
  • We conduct comprehensive experiments on both closed-source models (i.e. GPT-3.5, GPT-4 and OpenAI-o1) and open-source models (i.e. DeepSeek-V3, Deepseek-R1, and Llama-3.1). The results demonstrate that MigGPT achieved an average migration accuracy of 74.07%($\uparrow 45.92\%$), representing a significant improvement over vanilla LLMs.
@@ 2.1 Code Migration Kernel Patch
Existing code migration pycraft ; PPatHF technologies primarily focus on API cross-version maintenance diff_catchup ; a3 ; app_evovle ; cocci_evolve ; neat ; T2R ; REFAZER ; PYEVOLVE and the backporting of CVE security patches SKYPORT ; FixMorph ; PatchWeave ; TSBPORT . The most similar migration efforts, FixMorph FixMorph , TSBPORT TSBPORT , and PPatHF PPatHF , focus on CVE patch or forked code, which cannot be applied to the migration of out-of-tree kernel patch: 1) these methods only partially address out-of-tree kernel patch migration due to the tight coupling between kernel and patch code. They identify vulnerability patterns and apply predefined rules static_program_analysis but fail to manage complex changes such as namespace modifications, invocation conflicts, and control/data flow dependencies, limiting their effectiveness in comprehensive migration scenarios. 2) These works handle single-step migration, assuming known target code locations in updated repositories, making them difficult to apply to out-of-tree kernel patch scenarios. In contrast, our MigGPT tackles a more complex two-step task: retrieving target code from updated kernels and migrating patches, which is both more challenging and more practical.
@@ 2.2 LLMs for Coding
In recent years, LLMs codex ; incoder ; codellama ; coderl ; codegen ; starcoder ; gpt4 have achieved remarkable progress in various natural language processing tasks. Initially focused on natural language understanding and generation, the adaptability of LLMs has expanded to the field of software engineering, where they can be fine-tuned to perform programming tasks such as code completion rf-1 ; rf-2 ; rf-3 , code search codebert , code summarization rf-4 , code generation rf-5 , and even complex code repair vulrepair ; repair . This inspires us to apply LLMs to the migration of out-of-tree kernel patches. To the best of our knowledge, MigGPT is the first work to apply LLMs to this task, paving the way for subsequent research.
3 Problem Formulation
[…]
@@ 4.1 Migration Types
Table 1: Formalization and counts of the two types of migration examples. Other cases are too simple to necessitate resolution. Details in App. A.2.
Class Formalization Number
Type 1 $\Delta\neq\varnothing,\ \Sigma\neq\varnothing,$ 80
$\forall\delta\in\Delta,\ \forall\sigma\in\Sigma,\ \langle\delta,\sigma\rangle=0$ (59.3%)
Type 2 $\Delta\neq\varnothing,\ \Sigma\neq\varnothing,$ 55
$\forall\delta\in\Delta,\ \forall\sigma\in\Sigma,\ \langle\delta,\sigma\rangle\neq 0$ (40.7%)
Others $\Delta=\varnothing$ or Too simple
$\Delta\neq\varnothing,\ \Sigma=\varnothing$ to resolve
@@
Type 1: This type of migration example satisfies $\Delta\neq\varnothing,\Sigma\neq\varnothing$, $\forall\delta\in\Delta,\forall\sigma\in\Sigma,\langle\delta,\sigma\rangle=0$. This indicates that both the out-of-tree kernel patch and the new version of the Linux kernel have modified the code snippet, and their changes do not affect the same lines of code, meaning the modifications do not overlap or conflict with each other. Traditional methods’ limitations with Type 1 are discussed in App. G.1.
@@ 4.2 Benchmark Design
We have built a robust migration testing benchmark using out-of-tree kernel patches from real-world projects, specifically focusing on three open-source initiatives: RT-PREEMPT rt-preempt , Raspberry Pi Linux raspberrypi-linux , and HAOC haoc-patch   [^1] Even with knowledge of the code in these out-of-tree kernel patches, LLMs still struggle to accomplish migration and maintenance tasks.. More details about these projects are available in App. A.1. We collect code from these projects across Linux kernel versions 4.19, 5.4, 5.10, and 6.6 for our benchmark. The selected kernel versions are officially maintained LTS releases, which are widely adopted in production systems(e.g., enterprise servers, embedded devices).
@@ 5.1 Challenges
Challenge 1 (Structural Ambiguity): When identifying the code snippet $s_{\text{new}}$ in $\text{file}_{\text{new}}$ for migration, retrieval errors can occur. LLMs often struggle to locate function definitions within $s_{\text{new}}$ due to interference from similar function structures, leading to inaccuracies that affect subsequent migration stages. An example is provided in App. C.1.
Challenge 2 (Boundary Indeterminacy): This challenge occurs when retrieving $s_{\text{new}}$ from $\text{file}_{\text{new}}$. Due to the inherent randomness in LLM-generated responses, discrepancies often arise between the start and end lines of $s_{\text{new}}$ identified by the LLM and those retrieved by human developers ($\hat{s}_{\text{new}}$). This indeterminacy can result in missing or extraneous lines, significantly compromising migration outcomes where precise code segment boundaries are critical. An example is provided in App. C.2.
@@ 5.3 Code Fingerprint
Figure 3: A code snippet containing inline assembly statements and comment annotations.
To address the challenges LLMs face in migrating out-of-tree kernel patches across Linux kernel versions, a detailed analysis of code snippet structure is essential to identify migration-related code structure information and code scope constraints. While tools like Abstract Syntax Tree (AST) are useful for structural analysis, they have limitations: 1) Inability to process code snippets that lack complete compilation dependencies (e.g., missing variable or function definitions, absent macro declarations, or incomplete header inclusions) due to tight integration with the compilation process. 2) The mismatch between excessive structural details (AST tools provide a plethora of information irrelevant to patch migration) and the absence of critical information (such as comments and inline assembly), which is essential for maintaining and updating kernel patches  [^2] Inline assembly is widely used in the Linux kernel, and comments are crucial for future module development, as their omission would hinder subsequent modifications.. Focusing on key statements, such as migration points and alignment positions, while preserving essential elements like comments and inline assembly, can enhance efficiency and reduce overhead in the migration process.
@@
To address the limitations of traditional code structure analysis, we propose Code Fingerprint (CFP), a lightweight sequential data structure for analyzing code snippets. CFP records both the content and positional information for each line of statements, encompassing all C language statements, including comments and inline assembly (a detailed example is provided in App. D.1). As shown in Figure 4, CFP focuses on recording function definitions and function calls, which are crucial for addressing challenges 1 and 3, as detailed in Section 5.4. Additionally, its linear structure facilitates accurate positioning for insertion, deletion, and other update operations, tackling challenges 2 and 4, further explained in Sections 5.5 and  5.6. The algorithm for generating CFP is in App. D.4. Overall, CFP offers three key advantages: 1) effective processing of incomplete code snippets, 2) preservation of critical information such as comments and inline assembly, which are vital for out-of-tree kernel patch migration, and 3) a streamlined design that focuses on essential statements, improving migration efficiency and reducing overhead. CFP is specifically designed for out-of-tree kernel patch code, addressing the challenges encountered in Linux kernel patches. More discussions about CFP are in App. G.2. By minimizing unnecessary processing while ensuring relevance, CFP provides a targeted solution for migrating out-of-tree kernel patches across Linux kernel versions.